OFF RESONANCE FOUR WAVE MIXING STUDIES OF GROUND STATE MOLECULAR STRUCTURE FROM WEAK TO STRONG FIELDS

\underline{MATTHEW COMSTOCK}, IGOR PASTIRK, VADIM V. LOZOVOY, and MARCOS DANTUS, Michigan State University, East Lansing, MI, 48824.

With off resonance Four Wave Mixing (FWM) we can obtain background free rotational and vibrational information in the time domain. This information can be used to determine molecuar structure. We use this technique to study the effects of strong fields on molecular structure.